Your neighbor had his water analyzed because he was worried about lead in his water. His results said he had 10 ppm lead, 50 ppm radon, and 100 ppm iron in his water. If he had 1.00 kg sample of his tap water, how many grams of each contaminant is present?
1
Expert's answer
2017-11-29T13:01:06-0500
The “ppm” abbreviation stands for “part per million”, so 1 ppm means 1/1 000 000 part of the whole.
So if we have 1.00 kg of water than 1 ppm will be 1.00 kg / 1 000 000 = 0.001 g of contaminant.
Thus in case of 10 ppm of lead our sample contains 10 * 0.001 g = 0.01 g of lead; 50 ppm radon – 50 * 0.001 g = 0.05 g of radon; 100 ppm iron – 100 * 0.001 g = 0.1 g of iron.

Answer:
The sample contains 0.01 g of lead, 0.05 g of radon, 0.1 g of iron.